OpenStudy (anonymous):

(a+b)^2=a^2+b^2 true or false?

OpenStudy (anonymous):

@Hero @triciaal

OpenStudy (xapproachesinfinity):

what do you think try expanding (a+b)^2

OpenStudy (xapproachesinfinity):

(a+b)^2=(a+b)(a+b)

OpenStudy (xapproachesinfinity):

nope

OpenStudy (anonymous):

\[(a+b)(a+b)=a ^{2}+ab+ba+b ^{2}\]

OpenStudy (xapproachesinfinity):

correct! you need those middle terms why did you delete you reply lol

OpenStudy (triciaal):

false see the post by @Audeezy

OpenStudy (anonymous):

I got it, Thanks :)

OpenStudy (xapproachesinfinity):

W!